### v1

# SolNet: Open-source deep learning models for photovoltaic power forecasting across the globe

Abstract
Deep learning models have gained prominence in the field of solar forecasting. One drawback of these models is that they require a lot of high-quality data. This is often not feasible, especially in solar power forecasting due to poor measurement infrastructure in legacy systems and the rapid build-up of new systems. This paper proposes SolNet: a novel, general-purpose solar power forecaster, which addresses these challenges by using transfer learning from abundant synthetic data generated from PVGIS. Using actual production data from hundreds of sites in Australia, Belgium and the Netherlands, we show that SolNet improves forecasting performance over data-scarce settings as well as simple benchmark models. We observe this effect to be the strongest when only limited data is available, and show that seasonal patterns have a major impact on the results. The study also quantifies synthetic data requirements to achieve these improvements, as well as the effect of misspecifications.
keywords:
Energy forecasting , Transfer learning , Neural Networks , Synthetic data , Seasonality
journal: International Journal of Forecasting
1 Introduction
Renewable energy sources (RES) are seen as a critical component of the energy transition necessary for climate change mitigation (panwar2011role). However, the actual generation from these RES depends on local weather conditions and is known for its variability and intermittency. As such, expanding generation through RES introduces additional uncertainty in the power system operation, necessitating sophisticated generation forecasts on hour-ahead to day-ahead levels (notton2018intermittent). However, with rapidly growing solar and wind installed capacity, both the average and worst case forecast errors have also rapidly grown. Solar power forecasting is especially problematic, with the worst-case day-ahead generation forecast error (mean absolute error calculated on hourly forecasts) having grown from around 3 GW to over 13 GW over 16 European countries (kazmi2022good). Moreover, the proliferation of rooftop systems in the distribution grid further complicates grid operation by introducing over-voltage and reverse power flow concerns (steffel2012integrating). Furthermore, unlike large solar plants where information is often freely available, forecasting power generation from small scale distributed solar systems is complicated by the lack of supporting forecasts for relevant exogenous factors (e.g. ambient conditions, cloud cover, etc.).
Motivated by these concerns, solar power forecasting has seen significant progress as a field in recent years (hong2020energy). Inspired by the state-of-the-art performance of deep learning in other domains such as computer vision and natural language processing, several researchers have also utilized these methods to build solar forecasters (kumari2021deep). The results of naively applying deep learning in forecasting, however, are mixed. Makridakis et al. (2018) talked about the seemingly under-performing pure deep learning models, compared to hybrid or pure statistical models after analysing the results of the M4 forecasting competition (makridakis2018m4), a key competition in the field of time series forecasting.
The Great Energy Predictor (GEP-III) competition, focused on forecasting building energy demand, resulted in similar conclusions about gradient boosting based methods outperforming deep learning methods (miller2020ashrae). A major reason for deep learning underperforming simpler models in many forecasting tasks is its requirement for enormous amounts of data to generalize well. On the contrary, conventional time-series models (e.g. those from the SARIMAX family), and simpler machine learning based models (e.g. those built on tree-based methods) can generalize relatively well from limited amounts of data. Consequently, they do not suffer from the fact that sufficient data on solar operation or exogenous factors is often not readily available in practice.
1.1 State-of-the-art
In recent years, transfer learning has emerged as a key ingredient to improve the sample complexity of deep learning methods (i.e. data required to generalise well) (weiss2016survey) in several domains including computer vision (CV) (gupta2022deep), natural language processing (NLP) (ruder-etal-2019-transfer), and reinforcement learning tasks (zhu2023transfer). In this scheme, a source model is trained initially on a large corpus of data (derived from a known source domain), which is then fine-tuned using sparse or limited observational data (derived from the target domain of interest). When building deep learning models from scratch, the initial pre-training phase (utilizing data from a source domain) is missing, leading to poor generalisation due to data sparsity. In practice, the closer the source domain is to the target domain, the better the results of transfer learning.
One such recent example of transfer learning applied to PV power forecasting used a 3-layer LSTM architecture to predict PV power output, where the entire network was first trained on the source domain (solar irradiance), with the last layer fine-tuned on the target data (PV output) (zhou2020transfer). Interestingly, even though the source and target domain are somewhat different, the authors report improved predictive accuracy. Even more recently, abubakr2022unleashing have used transfer learning to predict solar irradiance at solar farms with limited data by using the abundant data available at one site. The model architecture was a Recurrent Neural Network (RNN), using historical solar irradiance data with up to four exogenous weather variables. This results in (slightly) better performance when using transfer learning compared to training the model directly on the limited data. The relatively small improvement in performance could well be due to the fact that the prediction time horizon is only one hour in the future.
These studies are emblematic of the issues hampering adoption of transfer learning in the energy field. On the one hand, transferring from data observed at other, similar sites is not fully generalisable, i.e. it introduces a new requirement during the model’s training process. zhang2018generative used a generative adversarial network (GAN) to create synthetic time series of solar generation, which could theoretically improve forecasting accuracy. However, studies like this one are not common in literature to the best of our knowledge. They also suffer from some of the same issues: the GAN (or other trained generative model) requires previously collected datasets, which is the actual issue to begin with.
The studies published so far also suffer from several limitations in terms of their evaluation robustness. They typically utilise very limited testing or future look-ahead periods, raising question marks about the statistical significance of the reported results. The seasonal variations of solar production over the year limits the generalisability of these results even more. Furthermore, they typically report improvements only in asymptotic performance, i.e. do not consider initial performance and learning performance. A more complete evaluation of transfer learning would look at differing amounts of training data, up until the initial performance, with no target data to fine-tune on. We elaborate on this further in the methodology section.
Finally, arguably the most important reason behind the wide-spread adoption of source models in computer vision and natural language processing has been their availability to researchers and practitioners alike. The few source models that have been trained by researchers for energy or solar applications are all invariably closed-source and inaccessible, having been trained from scratch themselves (HU201683; QURESHI2017742; zhou2020transfer; abubakr2022unleashing). This largely defeats the purpose of transfer learning.
1.2 Contributions
Consequently, in this paper, we propose a novel way to forecast residential PV using deep-learning: SolNet, a unique open-source deep learning model generator for solar power forecasting for any land-based site on the planet. The proposed models, made available for use through a Python package (GitHub link in the conclusion section), rely only on API access to PVGIS (HULD20121803), to generate PV production data for several variations of weather conditions and panel specifications (e.g. tilt, azimuth, etc.). This removes the need for data collection for end-users, and makes the transfer approach more realistic. Furthermore, the models can be fine-tuned with observational data where it exists.
We compare the performance of SolNet to benchmarks using data for several years from 40 test locations in the Netherlands, 299 in Australia and 2 (commercial sites) in Belgium, respectively. We report all three performance metrics for better understanding of the value of transfer: initial performance, learning improvement, and asymptotic performance. Figure 1 visualises this idea. The initial performance of our models gives us the error when we have no (target) data to train on. For the target model this means that we use randomly initialised weights. For the transfer model this means that the model has pre-initialized weights based on the training that happened in the source domain. The learning performance is the increased performance by adding more data to the training process and the asymptotic performance is the difference in performance between the target and transfer model that we expect to have over time when enough data is added.
Figure 1: Complete assessment of transfer performance
1.3 Organization
The remainder of this paper is structured as follows: Section 2 discusses the transfer learning methodology and the models utilised for forecasting. Section 3 elaborates on the data used in this transfer learning process, including the source (PVGIS) and target (the actual site locations) domains, and how they are prepared for the analyses. Section 4 provides detailed results from applying the proposed methodology, while Sections 5 and 6 conclude the paper, going over the most important findings as well as some future research directions.
2 Methodology
We first discuss the design of the forecasting exercise, as well as the benchmark used for comparison purposes of our transfer approach. Next, we discuss the architecture of our chosen neural network, LSTM (long short term memory), which is a type of recurrent neural network which can learn temporal associations. Next, we elaborate on the transfer learning methodology applied to the LSTM models. Finally, we talk about the metrics used for the evaluation.
2.1 The forecasting exercise and baseline approach
For this paper we focused on the case of forecasting the next 24 hours, from 00:00AM to 23:00PM with no gap between inference time and the forecast period. This is a simple reference case to showcase the performance of SolNet, however the functionality is available for users to change the point of inference as well as the forecast period and it is our intention to benchmark the performance using different windows in later research.
To compare the performance of our deep learning model, we use a Naive seasonal model (K=24). This model takes the corresponding hour of the previous day as the prediction for the next day. So, if the actual value for 14:00 at day t was 0.45 kWh, the value for 14:00 at t+1 will also be 0.45 kWh. This method provides us with a realistic baseline, given the diurnal pattern of solar power. We will refer to this method as the ’naive’ or ’naive seasonal’ method for the remainder of this paper.
2.2 LSTM as a forecast model
The deep learning architecture used in our experiments, both with and without transfer learning, is the Long Short-term Memory (LSTM) model (hochreiter1997long). An LSTM model consists of several identical units. These units have several gates that remember (or forget) information over time. This allows the model to retain a form of memory, making it possible to learn long term patterns in time series data. All these gates are made from the fundamental building blocks of neural networks: weights, biases and non-linearities in the form of activation functions (e.g. sigmoid or tanh functions). These weights and biases are the parameters that are learned during the training process. The below formulas provide an overview of the internals of the LSTM unit.
Input Gate: $i_{t}=\sigma(W_{xi}x_{t}+W_{hi}h_{t-1}+b_{i1})\times\tanh(W_{xi}x_{t}+W_{hi}h_{t-1}+b_{i2})$ (1)
Forget Gate: $f_{t}=\sigma(W_{xf}x_{t}+W_{hf}h_{t-1}+b_{f})$ (2)
Output Gate: $o_{t}=\sigma(W_{xo}x_{t}+W_{ho}h_{t-1}+b_{o})$ (3)
Cell State: $c_{t}=f_{t}c_{t-1}+i_{t}$ (4)
Hidden State: $h_{t}=o_{t}\tanh(c_{t})$ (5)
where:
$W_{xn}$ : are the weights for the input
$W_{hn}$ : are the weights for the hidden state
$b_{n}$ : are the biases
$h_{t-1}$ : is the hidden state at time t-1
$c_{t-1}$ : is the cell state at time t-1
Key to the LSTM are the cell state (the long-term memory) and the hidden state (the short-term memory). These two states get updated based on the states at the previous time step and the input associated with the current time step, after they pass through several gates. These gates manage what gets passed on to the next states (and what gets output).
The relative simplicity of the LSTM model, existing out of N layers with identical building blocks, taking (scaled) inputs and running them through several activation functions after applying weights and biases to them, makes it a good candidate for transfer learning. The fine-tuning options are (relatively) limited compared to some modern architectures, and the only prerequisite of using the model in both the target domain and the source domain, is that the inputs to the models are of equal shapes (i.e. without introducing additional complexity). In the univariate case, this prerequisite is easily fulfilled as the model only requires the user to have K hours of historical PV output data available, with K being the number of historical lags used to make the forecast. For the purpose of this research we limit this to K=24. From trial and error we noted that adding more lags did not add significantly to training performance, whilst increasing the training time by a significant margin. Due to this research using over 300 test sites, this additional training time was not practical to account for.
The final architecture of the LSTM used in our experiments was found after doing a grid search for the best hyper-parameters (layers, neurons per layer, dropout percentage), and includes 3 layers, each consisting of 400 LSTM units, with one more layer consisting of 24 linear units. These 24 units form the outputs, one for each hour of the day. Other hyper-parameters include the optimizer used during the back-propagation process, we use ADAM based on existing literature, the learning rate, which is $10^{-4}$ in the base case and $10^{-6}$ in the fine-tuning case (see later in this section), a dropout rate to avoid overfitting (50% in our case), and a batch size (32). These parameters (learning rate, dropout rate and batch size) were chosen based on trial and error, as well as common useage in literature.
2.3 Transfer learning
With the deep learning architecture in place, it is possible to explore the effect of transfer learning on solar forecasting in a principled manner. More concretely, the LSTM can be pre-trained on source data, after which it can be fine-tuned using target data to achieve potential improvements over the target model. Transfer learning, formally defined as improving the learning of a target predictive function on the learning task, $\mathcal{F}_{T}(\cdot)$ in the target domain, $\mathcal{D}_{T}$, using knowledge of $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$ (the source domain and the source task) (pan2010survey). While $\mathcal{D}_{T}\neq\mathcal{D}_{S}$ and $\mathcal{T}_{T}\neq\mathcal{T}_{S}$, we do assume that, the more similar the two are, the easier the transfer learning task will be. In our case, we keep the source domain as closely related to the target domain as possible by mimicking the location and the PV set-up of our target domain data, in our source domain, but we also provide results for what happens when this condition does not hold (see section 4.4). Concretely, in addition to directly using the pre-trained source model for the target task of solar PV prediction, there are a few different algorithmic ways in which transfer learning can be applied to the LSTM networks:
  1. 1.
    Direct application. The pre-trained source model can be directly applied in the target domain without any adjustments to the model parameters.
  2. 2.
    Finetuning. The source LSTM can be fine-tuned to the target domain. We keep the architecture intact, and run several additional training iterations using the (limited) target domain data. We can avoid model overfitting on the limited target data by lowering the learning rate, thereby making sure that the model only makes slight adjustments to the previously trained weights. Another popular method to fine-tune deep neural networks is by ”freezing” the parameters of certain units, or entire layers (see e.g. 9070676). In this case, the optimiser can no longer update the parameters in these layers during the optimisation process.
For our experiments, we combine both finetuning strategies: we freeze the initial LSTM layer, and lower the learning rate of the subsequent layers by a factor of 100, when retraining on the target domain data. This ensures that the fine-tuning process does not completely over-write the learning from the source domain. We also look at direct application of the source models. This basically represents the case where we only evaluate the initial performance and no learning performance.
2.4 Model evaluation
Performance is measured by way of the RMSE, in accordance with (yang2020verification). yang2020verification state that the RMSE is favored over the MAE because it penalizes large errors more and because it is more popular. The first reason makes sense in solar forecasting since it is important to correctly predict the timing and size of the daily peak in output. The second reason is essentially a ’network effect’, where the value of using something increases when more people use it. This makes sense, since we want results to comparable over studies, so using the same error metric as most people use makes sense. In all cases, a full year of target data is kept aside for final evaluation. 80% of the remaining data is used for training and 20% is used to validate the model performance and pick the best performing iteration of the model.
The complete workflow of the experiments is presented in figure 2. The first column shows the workflow for a source trained model. The middle column shows the process for a target trained model, skipping the source step and training directly on the target domain instead of finetuning with a lower learning rate and frozen layers. The final column shows the naive seasonal model, which takes the values of the day before as the predictions, requiring no training.
Figure 2: The workflow of the experiments
3 Data
Two types of data are required for this research: (1) the target domain data which has been observed in physical sites of interest but is limited in quantity, and (2) the source domain data, which is available in ample amounts, but has actually been synthetically generated from a (similar) setup. First, we briefly discuss the target domain datasets, i.e. PV installations from the Netherlands, Australia and Belgium, with only around two years of data per site. Second, we discuss the synthetic source domain used for transfer learning. Finally, we discuss the input features used in the forecast model, as well as the necessary pre-processing steps.
3.1 Target domain
3.1.1 Residential buildings in the Netherlands
Our main target domain data comes from different solar PV systems installed in several locations, both at the residential and commercial scale. The main bulk of the data comes from 40 individual household PV systems spread across the Netherlands. The data comes at a temporal resolution of 15 minutes (which we resample to an hourly resolution in accordance with the source domain). Even though we have more data available from Australia, the Dutch dataset is seen as our main corpus because we have site-specific meta-data as well, including the tilt, the azimuth, the peak power, and the approximate location of the PV systems (zip code and city). From this latter, we derive an approximate latitude and longitude. The peak power of the systems varies between 1.2 kWp and 3.1 kWp, with most systems around 2.5 kWp.
Data quantity varies by location: for all 40 locations, the full year of 2020 is available. Most locations also have data for the entire year of 2019 (although in some cases, this can be as few as three months only). As mentioned, the full year of 2020 is kept aside for evaluation of the final models. This fits well with the core premise of the research that there is often only a (very) limited amount of data available in the target domain before forecasts need to be made.
3.1.2 Residential buildings in Australia
We also utilise solar data from 300 Australian homes with rooftop solar systems spanning the period of 1 July 2011 to 30 June 2013. The solar systems have a mean size of 1.68 kWp and are all located in New South Wales (specifically in zip codes between 20 and 23). No meta-data is available for these locations other than the size of the system and the zip code.
3.1.3 Commercial buildings in Belgium
We also consider PV production data from two commercial buildings in Genk, Belgium. Data is available for the full year of 2019 for one location, and for nine months in the other location (the entire year of 2020 is available for testing, as before). In this case, the exact location of the systems is well-known, but the tilt and azimuth of these systems are not well-specified, because individual panels at the same site are placed in a variety of configurations. The systems have a peak power of 275 kWp and 125 kWp.
3.2 Source domain
The source domain data is derived from PVGIS, which uses two methods for calculating PV power data; satellite imagery and climate reanalysis data. The satellite images originate from the METEOSAT geostationary satellites, covering Africa, Europe and parts of Asia. From these satellite images, irradiation at ground level is derived using algorithms developed by CM SAF, creating the SARAH-2 dataset. For locations that do not have satellite data, climate reanalysis data is used, produced by the European Centre for Medium-range Weather Forecast (ECMWF), creating the ERA-5 data set.
For this research we use data generated by PVGIS from 2005 to 2019, mimicking the parameters (location, tilt, azimuth, peak power) of the previously discussed solar PV systems (in case these parameters are available) in order to maximize the similarity between source and target domain.
This data is subsequently used to train the source models we use for transfer learning. These models are then fine-tuned using the sparse observational target data. Even though PVGIS provides data for 2020, we do not use this year while creating source models as this can potentially cause data leakage. Data leakage refers to including knowledge about the future (or data which would realistically not be available) in a model. In this case, data leakage would occur if we were to use the year 2020 in the creation of our model and, simultaneously, evaluate the model on 2020. Similarly, for Australia we only use data up until 2011.
3.3 Feature engineering
By only looking at lagged load values and not including any external covariates, we keep the model as general as possible, requiring no additional data from a potential user of the source models. However, we can use the fact that we are dealing with time series data to include additional cyclical variables. Given the diurnal and annual patterns of solar power generation, it makes sense to add variables that capture these patterns. This is done by doing a cyclical encoding of the hours and months, by creating a sine and cosine feature for each of them. Featurising the pattern in this manner has several advantages. First, the cyclicality between the periods is respected, which makes it preferable to numerical encoding. Second, this method only creates 2 additional variables for the pattern, which makes it preferable to one-hot encoding.
3.4 Data pre-processing
In order to properly train and evaluate the results of our models and avoid leakage, the data has to be split into several sets. The source data is split in a train and test set, while the target data is split in a train set, a validation set, and a separate test set. We do this three-way split to mimic a realistic scenario where a proprietor of the system (e.g. a forecaster) will only have limited data (the train and validation set). So we cannot validate our models during training on this separate test set. We can however use this test set (the full year of 2020 for BE and NL, and the full year spanning 1 July 2012 to 30 June 2013 for AUS) to do a final evaluation of the fully trained model. We split the remaining data in a train and validation set, at a proportion of 80/20.
To train neural networks, our data should also be scaled to a similar range to accelerate convergence of the parameter estimation process. Given that all our variables are non-negative the data is scaled to be between 0 and 1 using a simple min-max scaler (see equation (6)). The scaling is fitted and applied on the train set, and also applied on the validation set, using the fit of the train set, to avoid any data leakage.
$f_{\mathrm{scaled}}(x)=\frac{x-x_{\min}}{x_{\max}-x_{\min}}$ (6)
There is an important caveat to the scaling however. While it is fine to use this type of scaling on the source data, it is not as simple in the target domain if we want to apply transfer learning. The model trained on the source domain data has been scaled on 12 years of data. We can therefore reasonably assume that the out-of-sample data will not, on average, be much different from the in-sample data. However, when we naively apply a min-max scaler to the target domain data, we get data that is, at times, incompatible with the models created in the source domain. For instance, consider the case where we only have access to target data from October to December. In the Northern hemisphere, these months have the lowest PV output. So, while in reality these lagged feature values should scale to, for instance, 0.6, we scale them to 1, negatively impacting the source model performance. We tackle this issue by normalizing using the peak-power (the nominal power on the DC side in kWp) times a factor of one minus 14%, the typical percentage used by PVGIS to take into account system losses. This peak-power is known in advance for all the systems, avoiding any data leakage while still keeping the data in the correct scale for using the transfer models. Thus, the scaling is still applied as shown in equation (6), with the addition that $x_{\text{max}}$ for the target data is now:
[…]
4 Results
In this section, we present results of applying transfer learning to residential PV systems in the Netherlands and Australia, as well as the commercial PV systems in Belgium. Next, using the Dutch data, we take a closer look at the impact of time of the year on model performance. To do this, we change the training period to assess the seasonal impact. Finally, we also assess the impact of data quantity in the source domain, decreasing the amount of data available from 15 years to 3 years, and the impact of data quality by increasing the (real) distance between the source and target domain by using source data from the South of France instead of the Netherlands.
4.1 Forecasting residential PV systems
The results for 40 residential PV systems in the Netherlands are summarised in figure 3. On the x-axis we plot the performance based on the months of data available in the target domain. ”0m” is the case with no target data (cf. 1. Likewise, ”1m” means that one month of target domain data used for training/finetuning. The error over time, of the baseline model (a seasonal naive model taking the values of the 24 hours of the previous day to be the new prediction), is a straight line, which does not improve over time, as it is independent of the amount of data available. On the y-axis we plot the RMSE for each of these cases. The thick lines are the overall median [^1] The median values are shown instead of mean values because not all datasets have the same observation period, which greatly impacts the mean estimate. values and the shaded areas represent the range of the individual cases. In other words, the shaded area represents the results of all 40 PV systems, from lowest to highest RMSE.
Figure 3: RMSE given months of target training data (NL)
From this plot, we can make four observations. First, the transfer models strictly outperform the baseline models. Second, the target model only starts to outperform the simple baseline once a sufficient amount of data (approx. 4 to 6 months of data depending on the test site) is available to train the model. This is very likely driven by the fact that this time period is sufficient for the model to have seen months with (near) optimal PV output. Third, we can see that the asymptotic performance improvement is fairly limited once enough data becomes available. The transfer model performance only improves by about 3.3% over the months, from 0.115 RMSE to 0.111. Fourth, the transfer model almost always outperforms the target model, and continues to do so even after most of the yearly seasonality is accounted for. Figure 4 shows us, in more detail, that the transfer model outperforms the target model in 418 out of 421 scenarios. This scatter plot also shows that most of the value from transfer learning lies in the first few months when insufficient target data is available, although for some systems performance improvements can be seen even after 10 months of target data are available.
Figure 4: Delta between RMSEs of transfer and target models (more negative is in favor of the transfer model)
Figure 5 plots the corresponding results for Australia. Given the larger size of the dataset (299 samples), the shaded area shows the 0.05 to 0.95 percentiles for each month to filter out some of the noise in the outliers. In a similar fashion to the Dutch results, most of the value lies in the initial months. As before, it takes over 7 months for the target model to consistently outperform the naive model here. It is worth noting that our target data to train on now ends in June, going backwards for increasing period size. Given that this is the winter period in Australia, the seasonality should not play a major role in the comparison of these two locations. The learning performance for the transfer model is again small. The median RMSE is 0.135 for the model with no fine-tuning (0m), improving by about 7.4% to 0.125 after 12 months. The target models outperform the transfer models in 8.6% of cases, mostly in the final four months.
Figure 5: RMSE given months of target training data (AUS)
4.2 Forecasting commercial PV systems
Figure 6 plots the results for the two office buildings in Belgium, showing that results are in line with the results from the Netherlands. The transfer model outperforms the target model, with most of the performance difference in the first six months. The forecasting accuracy, in general, is better compared to the Netherlands. This is clear from the baseline RMSEs, which are 0.108 and 0.116 compared to the median baseline of 0.141 in the Netherlands. This is most likely due to the larger scale of the PV installation on these office buildings, as well as the diverse angles of the panels. This causes the output to be smoother compared to smaller installations with one specific angle. This in turn makes the forecasting exercise easier.
Figure 6: RMSE given months of target training data (BE)
4.3 The effect of seasonality
Next, we focus on the aspect of yearly seasonality. Given that the target models catch up to the transfer learning performance after over half a year, it is reasonable to expect that the learner requires some of the summer (or transition) months to be trained on to achieve decent performance. The training so far was done chronologically with December as the final month, i.e. for the case of ’3m’ or three months of data, the data comes from the months of October, November and December (for NL and BE). The training and validation split was also done chronologically, i.e. the validation data for this ’3m’ case consists of the last 18 days of December.
To explore this further, we shifted the terminal month in time for different models. Table 1 shows these results for NL, by focusing specifically on what happens with the RMSE of the target models relative to the daily persistence model. The columns represent the final month of the target data used for training. The earlier in the year we have the terminal month, the less data we have in total, hence the increasing amount of n.a. values. We can see that there is a clear trend: the earlier in the year the final month is, the better. This most likely has to do with the presence of the best performing day inside the training data. Out of the 40 datasets, the month with the highest PV-output hour was in April 21 times and in May eleven times. As soon as this data is available to the model, the performance drastically increases. It is also the case that the main forecasting errors come from underestimating the peaks, which further explains the need for this data to be available during the training phase.
Table 1: Relative RMSE of target model compared to the baseline dependent on the terminal training month for the Dutch case
Dec Nov Oct Sep Aug Jul Jun May
0m 48% 47% 47% 47% 47% 47% 47% 47%
1m 37% 34% 26% 16% 14% 18% 25% 7%
2m 37% 27% 14% 30% 8% 5% 21% 2%
3m 26% 8% -5% -3% 6% -3% 8% -16%
4m 10% -12% -10% 2% 4% 15% -17% -16%
5m 4% -3% -12% 13% 3% -16% -20% n.a.
6m 6% -15% 0% 4% -17% -19% n.a. n.a.
7m -14% -17% -5% -14% -19% n.a. n.a. n.a.
8m -17% -17% -18% -18% n.a. n.a. n.a. n.a.
9m -22% -19% -21% n.a. n.a. n.a. n.a. n.a.
10m -21% -22% n.a. n.a. n.a. n.a. n.a. n.a.
11m -24% n.a. n.a. n.a. n.a. n.a. n.a. n.a.
4.4 The effect of source data quantity and quality
As a final step to test the generalisation of the model, we focused on the role of the source domain and the quality of the pre-trained model. The experimental setup proposed in this paper is unique in this aspect, as pre-training on prior observed data from other similar sites does not necessarily provide this flexibility. More specifically, we assess the impact of increasing the distance between the source and target domain by using source data from the south of France. Additionally, we also look at the impact of limiting the amount of data in the source domain to three years instead of fifteen. The key results can be seen in figure 7. With regards to the data quantity, we can see that the impact is very limited (only 0.002 in scaled RMSE). If we look at the impact of changing the source location (blue line), we see a more obvious effect. The initial performance even degrades to a worse performance than the baseline. Finetuning is clearly needed in this case, as the model has to adapt to the target location, and the learning performance is more substantial here. However, the performance appears to remain worse than a model trained on the ”correct” location with a difference in the scaled RMSE of 0.012 after almost an entire year.
These results further demonstrate that i) data quantity is only important up until the point where seasonality is fully accounted for and ii) the distance between the source and target domain plays a significant role in transfer performance, even though we still see initial improvements in mis-specified source models compared to target models.
Figure 7: The RMSE over time given a limited amount of data (orange) or data from a different location (blue)
5 Discussion
The findings in the result section are similar across all locations, with most of the performance gain offered by transfer learning, using the synthetic data, coming from the difference in initial performance. The learning performance is limited, and so is the asymptotic performance after a period of about six months. If data is available from a sunnier time period, this period of strong out-performance can be further shortened. This indicates that most of the difficulty is in correctly learning to forecast the seasonality inherent in PV energy production and, more importantly, the peak production. Additionally, it is important to try to mimic the situation of the target data (location, azimuth, slope) to limit the distance between source and target domain. However, the models do appear to be robust to small mis-specifications, and even exhibit improved accuracy for large mis-specifications.
5.1 Cumulative added-value
To further substantiate the significance of our research, we want to leave the reader with the following reasoning. While the asymptotic performance differential is weak between a target and transfer source model, having access to a performant model early on, when no (or limited) data is available, will have a large cumulative impact on the errors made. This can be seen in figure 8. In green we show the cumulative RMSE for the median transfer model, in blue for the median target model, and in orange for the baseline model. We can see that, over time, the impact of using the transfer model over the target model is quite large. The cumulative RMSE adds up to 1.354222Note that these RMSEs are on the zero to one scaled output for the transfer model compared to 1.954 for the target model, an improvement of over 30%. Having early access to such a model can also allow the PV owner or operator to identify commissioning or operational faults sooner.
Figure 8: The cumulative RMSE over time for different models
5.2 Limitations, and future directions
In this paper, we have considered the LSTM model as the neural architecture to train the solar power forecasters. While other, more recent deep learning architectures exist (e.g. transformer models, NBeats, Nhits) that can potentially achieve even better results, the goal of this paper was to provide a principled study into transfer learning as a tool to improve performance compared to simple statistical benchmarks on the one hand while simultaneously improving sample complexity of deep neural networks. We expect the results to generalize to these other neural architectures, but this is the subject of a follow-up study, similar to how successive GPT papers grew in complexity.
We have identified two key limitations of the paper. The first is the lack of access to datasets from climate zones outside of the temperate (maritime) climate zones of the Netherlands, Belgium, and even Australia, where all datasets originate from New South Wales. To further substantiate the added value of SolNet, the results shown in this study should be replicated in other parts of the world. We expect that, by making the code used for this paper freely available, more datasets will be used to validate the global performance of the transfer models. This will also allow other researchers to build on top of the models proposed here to achieve the same multiplier effect as seen in other domains such as computer vision etc.
A second limitation in the research is the lack of future covariates. The current paper exclusively used lagged PV output and cyclic variables to provide the most generalisable model possible, but adding forecasts of weather phenomena to the learning process (for example, forecasts of temperature, cloudiness or humidity) will further increase the performance of the models. Historical weather forecasts are more difficult to find however, certainly given that they have to be tailored to the specific timing and forecast horizon of the model. Together, tackling these two limitations, and integration of more deep learning models, are to be the subject of an extended, future SolNet.
6 Conclusion
This paper provides the first detailed assessment of the impact of utilising a large corpus of synthetic data in the creation of data-driven forecast models for solar power output. The improvements demonstrated hold regardless of the location of the PV panels, and are rather robust to mis-specifications in system parameters even. As evidence, the paper looked at over 300 sites in three different countries to substantiate the findings while running extensive tests lasting one year on all sites. This methodology extends existing recent work on transfer learning by minimising the need for data from ’similar sites’, and shows that there is a large boost to the initial predictive performance. As expected, this differential subsides over time, and is almost zero after a one year period, but the impact of using transfer models is accumulative over time, giving users a powerful tool to improve their forecasting performance at the early onset of (or even before) installing a new PV setup. The code, along with a functioning example for using SolNet for solar forecasting anywhere on land around the globe, has been open-sourced [^3] https://anonymous.4open.science/r/SolNet-CD07/README.md. We expect this work to be the first of many that have the potential to lead to general-purpose deep learning models for energy forecasting.